# Towards Aligning Multimodal LLMs with Human Experts: A Focus on Parent–Child Interaction

### v1

@@ Abstract.
Figure 1. Conceptual overview: We investigate how MLLMs can be guided to align with SLPs in parent–child interaction through two key stages—Observation, using structured prompts based on behavioural cues such as gaze, action, and vocalisation; and Judgement, informed by few-shot learning with expert-derived practices.:
@@ 1. Introduction
Early parent-child interactions lay the foundation for lifelong communication, social, and cognitive development. Foundational social-communicative behaviours–such as joint attention, joint intention, and social referencing–emerge during play and serve as critical indicators of developmental progress. Yet these behaviours are often subtle and difficult for caregivers to identify without professional support. This challenge is consequential: approximately one in six children in the United States experiences at least one developmental delay (robinson2017cdc), with language delays among the most prevalent (sunderajan_speech_2019). At the same time, access to speech-language pathologists (SLPs)–professionals who assess children’s communicative development by attending to multimodal cues such as gaze, gesture, vocalisation, and context–remains limited, particularly in rural or underserved communities (o2005barriers). As a result, expert developmental knowledge and care often remain out of reach in everyday caregiving contexts.
Despite recent advances in artificial intelligence (AI) for family and education (yuan2024designing; sun2024exploring; fiani2024exploring; nikkhah2024family; currin2024opportunities; su2024hidden), existing systems primarily assist with creating therapy materials or scaffolding caregiver engagement, rather than emulating the observing and judging practices of SLPs themselves. For example, Lewis et al. (lewis2025exploring) investigated how SLPs interact with AI-generated materials for children from culturally and linguistically diverse backgrounds, and Dangol et al.(dangol2025want) investigated how AI might support parents in delivering therapy at home, highlighting emotional and logistical challenges. Yet in both cases, the evaluative expertise of SLPs remained external to the AI systems. Little is known about whether MLLMs can be aligned to interpret parent-child interactions in ways that reflect expert reasoning.
In this paper, we explore how multimodal large language models (MLLMs) can be guided to emulate SLPs’ observing and judging practices. We focus on joint attention as a representative case: a core developmental milestone that SLPs typically evaluate through holistic, experience-based interpretations of children’s gaze, actions, and vocalisations. While MLLMs excel at structured tasks such as video captioning or temporal grounding, they struggle with socially nuanced behaviours that lack clear ground truth. We therefore investigate how expert criteria can be translated into structured representations that allow MLLMs to more closely align with expert judgment.
@@
  • We design and evaluate an exploratory MLLM system that aligns with speech-language pathologists’ approaches to joint attention assessment in two stages: (1) observing fine-grained behavioural cues from parent-child interaction videos using expert-informed prompting, achieving up to 85% accuracy across dimensions; and (2) evaluating interaction quality using only structured behavioural descriptions, reaching over 57% average accuracy compared to expert labels.
@@ 2.1. Language Delay and the Limits of Current Parent Support
Language delay is one of the most common developmental concerns in early childhood, with prevalence estimates ranging from 2.3% to 19% among children aged 2 to 7 years (mclaughlin2011speech). Without timely intervention, children with language delay are at increased risk for later difficulties in reading, attention regulation, and social interaction (sunderajan2019speech). SLPs play a key role in identifying and supporting children with language delays through structured interventions. However, limited access to professional services–whether due to long wait times, high cost, or service shortages–has led to growing interest in empowering parents to support their children at home (lieneman2017parent; straitstimes2024earlyintervention).
Parent training programmes such as It Takes Two to Talk, More Than Words, and DIR Floortime have shown success in encouraging behaviours that promote early communication and social development (pepper2004talk; sussman1999words; brock2014statewide; dir_floortime). These interventions typically guide caregivers to be more responsive and attuned during everyday interactions, and have demonstrated benefits in improving the quantity and quality of parent-child communication. Yet while such programmes can teach parents strategies to encourage communication, they rarely equip caregivers to evaluate the developmental effectiveness of those interactions (finke2009all; gadberry2011survey; ganz2013impacts).
For many parents, identifying nuanced developmental milestones can be difficult without professional guidance. This is especially true for joint attention (joint attention), one of the earliest and most important indicators of social-communicative development (mundy2007individual; tomasello1986joint). Determining whether joint attention is strong, weak, or absent often requires the kind of interpretive expertise that SLPs develop over years of practice and is challenging to translate into formal rules (o2005barriers). Families outside of clinical environments rarely benefit from such expert judgment. Motivated by this gap, we undertook this exploratory study oh how MLLMs might be aligned with SLPs’ reasoning to assess joint attention in naturalistic parent-child interactions. By probing the possibilities and limitations of this approach, we seek to inform future efforts to design AI tools that can augment professional expertise and broaden access to developmental support.
@@ 2.2. Technological Support for Parent-Child Interaction
There has been a surge in interactive technologies designed to support early childhood development, particularly by enhancing the quality and frequency of parent-child interaction. Early systems focused on lightweight support through manual logging. For instance, Chan et al. (chan2017wakey) developed WAKEY, which aimed to smooth morning routines by prompting parents to record schedules and track phrase usage. While effective in structuring communication, WAKEY relied entirely on manual input, without the ability to capture richer contextual signals.
Subsequent work shifted towards real-time speech analysis to provide parents with more immediate feedback. TalkBetter (hwang2014talkbetter) analysed turn-taking and generated tailored recommendations to scaffold language development, while TalkLIME (song2016talklime) extended this approach by offering live metrics such as utterance count, initiation ratio, and turn-taking frequency.
More recent systems incorporated multimodal sensing to deepen their understanding of interaction dynamics. Captivate! (kwon2022captivate) leveraged gaze estimation and speech recognition to detect joint attention and suggest parent prompts during play, and AACessTalk (choi2025aacess) supported minimally verbal autistic children by combining vocabulary recommendations with contextual guidance, ultimately improving turn-taking and parental reflection over a two-week deployment.
Parallel to these system developments, researchers have also explored how parents can be supported in adopting therapy strategies more effectively. Dangol et al. (dangol2025want) investigated how AI might address emotional and logistical barriers to at-home therapy, while Li et al. (li2025asd) introduced ASD-HI, a multimodal dataset and baseline model designed to detect strategies and assess fidelity in parent-led interventions.
@@ 2.3. Technological Support for SLPs
Generative AI is beginning to influence many aspects of clinical practice, including how SLPs source or adapt therapy materials. For example, Lewis et al. (lewis2025exploring) examined how SLPs interact with AI-generated visual content when working with children from culturally and linguistically diverse backgrounds. While their study highlighted important concerns regarding representation, bias, and contextual mismatch, it focused primarily on content usability. Their work does not address how AI might support or simulate the evaluative reasoning processes that SLPs use when observing and interpreting children’s behaviours in real time. With the rapid progress of MLLMs across tasks such as video captioning and temporal grounding (lu2024gpt), researchers have begun to explore their potential for analysing human behaviour–particularly in socially grounded contexts such as parent-child interaction. MLLMs can interpret complex verbal and non-verbal behaviour across video, audio, and text, and generate descriptive summaries of human activity (gandhi2023multimodal; jain2024vcoder). Thanks to their in-context learning capabilities, MLLMs can flexibly adapt to new interaction settings with minimal data (dos2023composite; whitehead2024generative), making them promising for contexts that require domain-specific reasoning. Crucially, these models shift the output from fixed metrics or visualisation into natural language descriptions–potentially bridging the gap between raw behavioural signals and expert interpretation. Zheng et al.(zheng2024soap) introduced SOAP.AI, a system that enables experts to interact with MLLMs through in-context prompting and collaborative task design, and supports the automatic generation of initial behavioural segments–such as those relevant to joint attention–in domains like healthcare and education. However, SOAP.AI was designed primarily as a tool for expert users, without evaluating the reliability of its outputs or aligning them with SLP judgement. Extending this line of inquiry, Zheng et al.(zheng2025ai) subsequently explored how AI-generated documentation could more directly support SLPs, identifying four opportunities and proposing three fluidity-focused design goals for future systems. However, while SOAP.AI exemplifies how AI can augment expert workflows, it is designed primarily for expert users and does not evaluate the reliability of its outputs or align them with SLP judgement, leaving open the question of whether MLLMs can simulate expert evaluative capabilities.
@@
While these approaches centre on supporting SLPs in their clinical workflows, they do not address a critical question: can current technologies simulate the observational and interpretive reasoning that SLPs apply when evaluating parent-child interactions? Bridging this gap requires rethinking how AI systems can align with expert judgement–not just to assist SLPs, but to empower parents with meaningful, expert-level insights into their everyday interactions. Our work explores the foundational alignment of MLLMs with SLP reasoning, laying the basis for more reliable support to both experts and caregivers.
2.4. Towards SLP-Aligned Observation and Judgement in Parent-Child Interaction
The first attempt to align MLLMs with the evaluative judgement of SLPs was introduced by Shi et al. (shi2025towards), who prompted MLLMs with definitions of joint attention to directly identify strong or poor joint attention segments through a temporal grounding task. Their results showed poor performance, primarily due to the models’ limited sensitivity to fine-grained gaze cues–particularly eye contact, which is central to joint attention dynamics. These findings highlight the need for closer alignment between MLLM interpretation and the evaluative reasoning used by SLPs, rather than relying solely on definition-based prompting.
@@ 3.1. Dataset
Figure 2. Flow chart of the video screening process :
@@
The videos in the dataset covers three categories: (1) behavioural guidance and skill modelling [^1] https://www.youtube.com/watch?v=YUkujhg6j6w (n=10), where parents demonstrate techniques such as PRIDE skills (masse2018taking) or “Big Ignore” (woodfield2021time); (2) language and cognitive development [^2] https://www.youtube.com/watch?v=rVqJacvywAQ (n=9), showing tasks like topic discussions or Piaget’s experiments; and (3) daily life skills and interaction [^3] https://www.youtube.com/watch?v=N3wAPLXd7I0 (n=6), including reward chart reviews or shared playtime.
@@
(a) Behavioural Guidance and Skill Modelling: Parent uses the ”Big Ignore” (woodfield2021time) technique while the child continues painting (Video 3 (lieneman2023)).
(b) Language and Cognitive Development: Parent and child play a toy-hiding game to foster engagement. (Video1 10 (Lieneman2024))
(c) Daily Life Skills and Interaction: Parent and child build a marble run together, encouraging planning (Video2 (adam2013)).
Figure 3. Examples from three categories in our dataset: behavioural guidance, language development, and daily life interaction.:
@@ 3.2. Participants
Our three expert SLPs were all female, aged between 30 and 59 years (see Table 1). All have extensive clinical experience working directly with children, particularly those on the autism spectrum, and have received training in a wide range of evidence-based intervention programmes, including Hanen’s It Takes Two to Talk (pepper2004talk), More Than Words and DIR Floortime (dir_floortime).
@@ 3.3. Study Protocol
Each SLP then reviewed all 25 videos. To focus our analysis and ensure task feasibility, we limited annotations to the child’s joint attention behaviour, which is often a primary concern in early developmental assessment. They marked the start and end times of segments they considered to show strong or poor joint attention, with all unlabelled segments treated as moderate by default. Each annotation included their short explanation justifying their decision and their suggestions for how Parent-Child Interaction could be supported. We used our custom-built video annotation tool to facilitate the SLPs’ judgement process (see Figure 4, shown here with a screenshot from Video 24444https://www.youtube.com/watch?v=rVqJacvywAQ (adam2013)).
@@
Figure 4. Our video annotation tool supports SLPs’ judgement process with four main components: Video Playback Area for watching, pausing, and replaying parent–child interactions (shown here with a screenshot from Video 24. Timeline Annotation Area for selecting and labelling segments as strong or poor joint attention; Note-Taking Area for recording justifications or observations; and Control Button Area for task submission and navigation. :
@@ 3.4. Data Analysis
The audio recordings were transcribed using WhisperX (bain2023whisperx) and manually verified. We analysed the qualitative data by combining the SLPs’ transcripts from the interviews with their annotations during the video tasks. To facilitate comparison across experts and to quantify areas of agreement and divergence, we standardised the data by dividing each video into fixed 5-second segments and mapping SLP labels onto these units. Any segments not explicitly labelled as Strong or Poor were automatically assigned a Moderate label.

### v2

@@ Abstract.
Figure 1. Conceptual overview: We investigate how MLLMs can be guided to align with SLPs in parent–child interaction through two key stages—Observation, using structured prompts based on behavioural cues such as gaze, action, and vocalisation; and Judgement, informed by few-shot learning with expert-derived practices.
This diagram illustrates how AI understands parent-child interactions by aligning with Speech-Language Pathologists (SLPs) and using cues like gaze, action, and vocalization.
@@ 1. Introduction
Early parent-child interactions lay the foundation for lifelong communication, social, and cognitive development. Foundational social-communicative behaviours–such as joint attention, joint intention, and social referencing–emerge during play and serve as critical indicators of developmental progress. Yet these behaviours are often subtle and difficult for caregivers to identify without professional support. This challenge is consequential: approximately one in six children in the United States experiences at least one developmental delay (Robinson, 2017), with language delays among the most prevalent (Sunderajan and Kanhere, 2019a). At the same time, access to speech-language pathologists (SLPs)–professionals who assess children’s communicative development by attending to multimodal cues such as gaze, gesture, vocalisation, and context–remains limited, particularly in rural or underserved communities (O’Callaghan et al., 2005). As a result, expert developmental knowledge and care often remain out of reach in everyday caregiving contexts.
Despite recent advances in artificial intelligence (AI) for family and education (Yuan et al., 2024; Sun et al., 2024; Fiani et al., 2024; Nikkhah et al., 2024; Currin et al., 2024; Su et al., 2024), existing systems primarily assist with creating therapy materials or scaffolding caregiver engagement, rather than emulating the observing and judging practices of SLPs themselves. For example, Lewis et al. (Lewis et al., 2025) investigated how SLPs interact with AI-generated materials for children from culturally and linguistically diverse backgrounds, and Dangol et al.(Dangol et al., 2025) investigated how AI might support parents in delivering therapy at home, highlighting emotional and logistical challenges. Yet in both cases, the evaluative expertise of SLPs remained external to the AI systems. Little is known about whether MLLMs can be aligned to interpret parent-child interactions in ways that reflect expert reasoning.
In this paper, we explore how multimodal large language models (MLLMs) can be guided to emulate SLPs’ observing and judging practices. We focus on joint attention as a representative case: a core developmental milestone that SLPs typically evaluate through holistic, experience-based interpretations of children’s gaze, actions, and vocalisations. While MLLMs excel at structured tasks such as video captioning or temporal grounding (Lu et al., 2024), they struggle with socially nuanced behaviours that lack clear ground truth. At the same time, recent advances suggest that MLLMs can interpret complex verbal and non-verbal behaviour across video, audio, and text, and generate descriptive summaries of human activity (Gandhi et al., 2023; Jain et al., 2024), and continues to be an active area of research. Their ability to adapt to task-specific criteria through in-context learning makes them a promising candidate for bridging raw behavioural signals with the structured interpretive practices used by SLPs. We therefore investigate how expert criteria can be translated into structured representations that allow MLLMs to more closely align with expert judgment.
@@
  • We design and evaluate an exploratory MLLM system that aligns with speech-language pathologists’ approaches to joint attention assessment in two stages: (1) observing fine-grained behavioural cues from parent-child interaction videos using expert-informed prompting, achieving up to 85% accuracy across dimensions; and (2) evaluating interaction quality using only structured behavioural descriptions, reaching over 64% average accuracy compared to expert labels.
@@ 2.1. Language Delay and the Limits of Current Parent Support
Language delay is one of the most common developmental concerns in early childhood, with prevalence estimates ranging from 2.3% to 19% among children aged 2 to 7 years (McLaughlin, 2011). Without timely intervention, children with language delay are at increased risk for later difficulties in reading, attention regulation, and social interaction (Sunderajan and Kanhere, 2019b). SLPs play a key role in identifying and supporting children with language delays through structured interventions. However, limited access to professional services–whether due to long wait times, high cost, or service shortages–has led to growing interest in empowering parents to support their children at home (Lieneman et al., 2017; The Straits Times, 2024).
Parent training programmes such as It Takes Two to Talk, More Than Words, and DIR Floortime have shown success in encouraging behaviours that promote early communication and social development (Pepper and Weitzman, 2004; Sussman, 1999; Brock et al., 2014; The Interdisciplinary Council on Development and Learning, 2025). These interventions typically guide caregivers to be more responsive and attuned during everyday interactions, and have demonstrated benefits in improving the quantity and quality of parent-child communication. Yet while such programmes can teach parents strategies to encourage communication, they rarely equip caregivers to evaluate the developmental effectiveness of those interactions (Finke et al., 2009; Gadberry, 2011; Ganz et al., 2013).
For many parents, identifying nuanced developmental milestones can be difficult without professional guidance. This is especially true for joint attention (joint attention), one of the earliest and most important indicators of social-communicative development (Mundy et al., 2007; Tomasello and Farrar, 1986). Determining whether joint attention is strong, weak, or absent often requires the kind of interpretive expertise that SLPs develop over years of practice and is challenging to translate into formal rules (O’Callaghan et al., 2005). Families outside of clinical environments rarely benefit from such expert judgment. Motivated by this gap, we undertook this exploratory study oh how MLLMs might be aligned with SLPs’ reasoning to assess joint attention in naturalistic parent-child interactions. By probing the possibilities and limitations of this approach, we seek to inform future efforts to design AI tools that can augment professional expertise and broaden access to developmental support.
@@ 2.2. Technological Support for Parent-Child Interaction
There has been a surge in interactive technologies designed to support early childhood development, particularly by enhancing the quality and frequency of parent-child interaction. Early systems focused on lightweight support through manual logging. For instance, Chan et al. (Chan et al., 2017) developed WAKEY, which aimed to smooth morning routines by prompting parents to record schedules and track phrase usage. While effective in structuring communication, WAKEY relied entirely on manual input, without the ability to capture richer contextual signals.
Subsequent work shifted towards real-time speech analysis to provide parents with more immediate feedback. TalkBetter (Hwang et al., 2014) analysed turn-taking and generated tailored recommendations to scaffold language development, while TalkLIME (Song et al., 2016) extended this approach by offering live metrics such as utterance count, initiation ratio, and turn-taking frequency.
More recent systems incorporated multimodal sensing to deepen their understanding of interaction dynamics. Captivate! (Kwon et al., 2022) leveraged gaze estimation and speech recognition to detect joint attention and suggest parent prompts during play, and AACessTalk (Choi et al., 2025) supported minimally verbal autistic children by combining vocabulary recommendations with contextual guidance, ultimately improving turn-taking and parental reflection over a two-week deployment.
Parallel to these system developments, researchers have also explored how parents can be supported in adopting therapy strategies more effectively. Dangol et al. (Dangol et al., 2025) investigated how AI might address emotional and logistical barriers to at-home therapy, while Li et al. (Li et al., 2025) introduced ASD-HI, a multimodal dataset and baseline model designed to detect strategies and assess fidelity in parent-led interventions.
@@ 2.3. Technological Support for SLPs
Generative AI is beginning to influence many aspects of clinical practice, including how SLPs source or adapt therapy materials. For example, Lewis et al. (Lewis et al., 2025) examined how SLPs interact with AI-generated visual content when working with children from culturally and linguistically diverse backgrounds. While their study highlighted important concerns regarding representation, bias, and contextual mismatch, it focused primarily on content usability. Their work does not address how AI might support or simulate the evaluative reasoning processes that SLPs use when observing and interpreting children’s behaviours in real time. With the rapid progress of MLLMs across tasks such as video captioning and temporal grounding (Lu et al., 2024), researchers have begun to explore their potential for analysing human behaviour–particularly in socially grounded contexts such as parent-child interaction. MLLMs can interpret complex verbal and non-verbal behaviour across video, audio, and text, and generate descriptive summaries of human activity (Gandhi et al., 2023; Jain et al., 2024). Thanks to their in-context learning capabilities, MLLMs can flexibly adapt to new interaction settings with minimal data (Dos Santos Melicio et al., 2023; Whitehead et al., 2024), making them promising for contexts that require domain-specific reasoning. Crucially, these models shift the output from fixed metrics or visualisation into natural language descriptions–potentially bridging the gap between raw behavioural signals and expert interpretation. Zheng et al.(Zheng et al., 2024) introduced SOAP.AI, a system that enables experts to interact with MLLMs through in-context prompting and collaborative task design, and supports the automatic generation of initial behavioural segments–such as those relevant to joint attention–in domains like healthcare and education. However, SOAP.AI was designed primarily as a tool for expert users, without evaluating the reliability of its outputs or aligning them with SLP judgement. Extending this line of inquiry, Zheng et al.(Zheng et al., 2025) subsequently explored how AI-generated documentation could more directly support SLPs, identifying four opportunities and proposing three fluidity-focused design goals for future systems. However, while SOAP.AI exemplifies how AI can augment expert workflows, it is designed primarily for expert users and does not evaluate the reliability of its outputs or align them with SLP judgement, leaving open the question of whether MLLMs can simulate expert evaluative capabilities.
@@
While these approaches centre on supporting SLPs in their clinical workflows, they do not address a critical question: can current technologies simulate the observational and interpretive judging that SLPs apply when evaluating parent-child interactions?
While general-purpose systems such as Personalized Judge (Dong et al., 2024) show that LLM-as-Judge can be effective in broad settings, their limitations become clear in expertise-heavy domains. Szymanski et al. (Szymanski et al., 2025; Chen et al., 2024) show that although LLMs align moderately with subject-matter experts in broad fields such as dietetics and mental health (64–68% agreement), their consistency drops on domain-specific questions that rely on professional judgement. Parent–child interaction illustrates this gap even more sharply: behaviours like joint attention require interpreting subtle, context-dependent cues rather than following fixed criteria. Prior attempts to prompt MLLMs for joint-attention grounding (Shi et al., 2025) yielded near-chance performance, largely due to the models’ insensitivity to fine-grained gaze signals. These patterns suggest that, despite general promise, LLM-as-Judge remains insufficient for specialised behavioural assessment without deeper alignment to expert reasoning.
@@ 3.1. Dataset
Figure 2. Flow chart of the video screening process
This diagram illustrates our video screening process, which began with a keyword search that yielded 198 videos. We then excluded videos based on our inclusion and exclusion criteria, resulting in a final set of 25 videos for the study.
@@
The videos in the dataset covers three categories: (1) behavioural guidance and skill modelling [^1] https://www.youtube.com/watch?v=YUkujhg6j6w (n=10), where parents demonstrate techniques such as PRIDE skills (Masse et al., 2018) or “Big Ignore” (Woodfield et al., 2021); (2) language and cognitive development [^2] https://www.youtube.com/watch?v=rVqJacvywAQ (n=9), showing tasks like topic discussions or Piaget’s experiments; and (3) daily life skills and interaction [^3] https://www.youtube.com/watch?v=N3wAPLXd7I0 (n=6), including reward chart reviews or shared playtime.
@@
(a) Behavioural Guidance and Skill Modelling: Parent uses the ”Big Ignore” (Woodfield et al., 2021) technique while the child continues painting (Video 3 (Lieneman, 2023a)).
(b) Language and Cognitive Development: Parent and child play a toy-hiding game to foster engagement. (Video1 10 (Lieneman, 2023b))
(c) Daily Life Skills and Interaction: Parent and child build a marble run together, encouraging planning (Video2 (Adam, 2013)).
Figure 3. Examples from three categories in our dataset: behavioural guidance, language development, and daily life interaction.
This diagram shows example videos from our dataset, categorized into three types: Behavioural Guidance and Skill Modelling (left), where a parent uses the ’Big Ignore’ technique while a child paints; Language and Cognitive Development (middle), where a parent and child play a toy-hiding game; and Daily Life Skills and Interaction (right), where they build a marble run together to encourage plannings
@@ 3.2. Participants
Our three expert SLPs were all female, aged between 30 and 59 years (see Table 1). All have extensive clinical experience working directly with children, particularly those on the autism spectrum, and have received training in a wide range of evidence-based intervention programmes, including Hanen’s It Takes Two to Talk (Pepper and Weitzman, 2004), More Than Words and DIR Floortime (The Interdisciplinary Council on Development and Learning, 2025).
@@ 3.3. Study Protocol
Each SLP then reviewed all 25 videos. To focus our analysis and ensure task feasibility, we limited annotations to the child’s joint attention behaviour, which is often a primary concern in early developmental assessment. They marked the start and end times of segments they considered to show strong or poor joint attention, with all unlabelled segments treated as moderate by default. Each annotation included their short explanation justifying their decision and their suggestions for how Parent-Child Interaction could be supported. We used our custom-built video annotation tool to facilitate the SLPs’ judgement process (see Figure 4, shown here with a screenshot from Video 24444https://www.youtube.com/watch?v=rVqJacvywAQ (Adam, 2013)).
@@
Figure 4. Our video annotation tool supports SLPs’ judgement process with four main components: Video Playback Area for watching, pausing, and replaying parent–child interactions (shown here with a screenshot from Video 24. Timeline Annotation Area for selecting and labelling segments as strong or poor joint attention; Note-Taking Area for recording justifications or observations; and Control Button Area for task submission and navigation.
This figure shows our video annotation tool, designed to support SLPs in their judgment process. It features four main components: a Video Playback Area (shown with a screenshot from Video 24) for watching interactions, a Timeline Annotation Area for labeling joint attention as ”strong” or ”poor,” a Note-Taking Area for recording observations, and a Control Button Area for task submission and navigation.
@@ 3.4. Data Analysis
The audio recordings were transcribed using WhisperX (Bain et al., 2023) and manually verified. We analysed the qualitative data by combining the SLPs’ transcripts from the interviews with their annotations during the video tasks. To facilitate comparison across experts and to quantify areas of agreement and divergence, we standardised the data by dividing each video into fixed 5-second segments and mapping SLP labels onto these units. Any segments not explicitly labelled as Strong or Poor were automatically assigned a Moderate label.